# Thickness dependent rare earth segregation in magnetron deposited NdCo4.6 thin films studied by Xray reflectivity and Hard Xray photoemission

### v3

@@ Abstract
Magnetic anisotropy in disordered rare earth–transition metal (RE–TM) compounds arises from RE atoms occupying asymmetric environments within the TM sublattice. However, the underlying mechanism that promotes such environments remains not fully understood. In this study, we investigate amorphous NdCo4.6 thin films deposited by magnetron sputtering, where the magnetic anisotropy evolves with thickness from in-plane to out-of-plane orientation above 40 nm. X-ray reflectivity measurements revealed the progressive formation of an additional layer between the 3 nm Si capping layer and the NdCo film with increasing film thickness. To probe the composition and distribution of cobalt and neodymium near the surface, Hard X-ray Photoemission Spectroscopy (HAXPES) was performed on films ranging from 5 nm to 65 nm in thickness, using incident photon energies of 7, 10, and 13 keV. These correspond to inelastic electron mean free paths of 7.2–12.3 nm in cobalt. The cobalt atomic concentration, deduced from HAXPES at the Nd 3d and Co 2p excitations, was consistently below the nominal value and varied with both thickness and photon energy. This indicates segregation of RE atoms at the film surface, which becomes more pronounced with increasing thickness. Background analysis of the Co 2p and Nd 3d peaks supports this conclusion. The thickness of the Nd segregated layer is estimated in 2-3 nm. These findings demonstrate that incorporating neodymium into the cobalt lattice incurs an energy cost which is associated with strain due to the volume mismatch between the two elements. Reducing this strain energy promotes anisotropic atomic environments for Nd, thereby explaining the emergence of perpendicular anisotropy and its dependence on film thickness in these RE-TM compounds.
@@ I INTRODUCTION
The PMA of RE-TM compounds is caused by the strong spin-orbit coupled 4f electron of the RE atoms. This orbital is nearly isolated from the valence atoms. Magnetic coupling between the spin of the TM and the 4f electrons occurs owing to the intratomic exchange with their valence 6s and 5d electrons which are spin polarized by the 3d TM electrons [7]. This coupling is weak, in the range of meV, and competes in energy with the crystal field at the RE[8, 7] which fixes the orientation of its 4f angular moment that defines the anisotropy of the compound. The asymmetry of the RE atomic environment is, therefore, key for understanding the PMA properties of RE-TM alloys[9, 10, 11]. Structural asymmetry has been demonstrated in other RE-TM compounds such as TbFe by difraction[11], transmission electron microscopy[12] and extended Xray absorption fine spectroscopy[13, 14].
Understanding how RE enters this type of asymmetric environment is an old problem owing to the complexity of the system[15, 16, 17]. However, the most accepted explanation for thin films deposited using magnetron sputtering methods, which have relatively low incident adatom energies, is based on lowering the surface energy during thin film growth[18, 19]. Following this explanation, a majority of RE are in asymmetric environments because the corresponding adatom locations during growth are those that reduce the surface energy the most, creating a structural texture that gives rise to the resulting anisotropic environments. Thus, to control the PMA and the magnetic properties of the NdCo compounds it is neccesary to elucidate the parameters that condition the adatom diffusion during thin film growth.
@@
The PMA properties of our NdCo4.6 are thickness dependent. PMA appears above a threshold thickness between 30 to 40 nm and gradually increases until it reaches its saturation value at a thickness below 100 nm, indicating that it is built during the growth process of the film. Therefore, surface-sensitive techniques applied to different thin-film thicknesses should provide information on how PMA is formed. For this study we used X-ray reflectivity (XRR) and HAXPES[29, 29]. XRR is highly sensitive to layer interfaces and should detect changes in thickness, electron density and roughness. However, it cannot provide the precise nature of the elemental composition of the layers and interfaces without ambiguity. The electron densities of cobalt and neodymium may not be large enough to be distinguished if there is some intermixing. Photoemission spectroscopy is an elemental-sensitive probe with probe depths that can be about three times the inelastic electron mean free path () of the photoemitted electrons[30]. By using hard Xrays, HAXPES reaches larger probing depths than Xray absorption spectroscopy, when using secondary electron detection [31], or Xray photoemission (XPS). Moreover, the probe depth can be modified by the incident photon energy and/or by choosing the specific excitation peaks[29]. This allows the study of the composition and chemistry of the top layer region of the samples underneath the 3 nm thick silicon capping layer and checks the extension of the possible segregated layers by changing the probe depth.
@@
The experiment is especially centered in the information collected by HAXPES to determine the gradation in the composition of thin films[32]. This information comes from the measured atomic relative concentrations of cobalt to neodymium, the valence band spectra, the shape of the photoemitted Co 2p and Nd 3d peaks and the shape and intensity of their background due to inelastically scattered electrons[33, 34]. The results of the experiment are consistent with an effective segregated layer of neodymium in all the thin films analyzed whose effective thickness is of the order of 2 to 3 nm, demonstrating that the incorporation of neodymium in cobalt is not an energetically favorable process, what explains the origin of PMA in the analyzed RE-TM compound based on the strain relaxation due to the atomic volumen difference between Nd and Co.
@@ II EXPERIMENT
The samples were co-deposited using three confocal magnetron sputtering guns, Co, Nd and Si. The Co gun was normal to the substrate whereas the Nd and Si guns formed an angle of 30∘ on opposite sides to the normal of the sample. The Ar pressure during deposition was 3 mbar, with a base pressure of 1 mbar. The thin films were deposited on the native oxide of boron doped silicon substrates at a rate of 2 Å. All the films had a protective capping layer of 3 nm thick silicon. Deposition rates and concentration were calibrated using a quartz balanced and XRR.
@@ III.1 Magnetic Characterization
Figure 1 shows the in-plane HLs obtained by magnetometry at RT for pure cobalt and NdCo4.6 thin films of similar thickness. The shape of the HLs of cobalt was independent of the film thickness. This is shown in figure 1(a) where the HLs of the cobalt thin films with thicknesses ranging from 5 to 50 nm are superimposed. Their coercive field is 20 Oe, which is practically the same as that measured in the NdCo4.6 thinner films (fig. 1(b)). Owing to the normal incidence of Co atoms, they have negligible in-plane magnetic anisotropy. Their magnetic remanence was close to 60.
@@
The HLs measured in samples with thicknesses below 40 nm are typical of soft magnets with in-plane anisotropy. Figure 1(b) shows them superimposed because they barely change with thickness. Their magnetic easy axis (EA) is perpendicular to the plane of incidence of the Nd and Si atoms, confirming the oblique incidence origin of their in-plane anisotropy[37, 38, 39]. Their EA HLs have almost 100 remanence, indicating that PMA energy is small compared with the in-plane anisotropy at these thicknesses. The hard axis (HA) has a saturation field, , of approximately 200 Oe at 30 nm, increasing to almost 400 Oe at lower thicknesses. The opening of the hysteresis loop in the HA might indicate a possible out of plane anisotropy component. As the thickness of the samples increased above 40 nm, the increased, the remanence decreased and the in-plane increased, as shown in figure 1(c). The HL of the 65 nm thick sample, in figure 1(d), shows a which is 10 times the measured in the samples with thickness lower than 40 nm. In addition, there is almost no difference between the HL measured along the in-plane EA and the HA: the in-plane anisotropy becomes overwhelmed by the out-of-plane anisotropy.
@@
The magnetization of both cobalt and NdCo4.6 thin films is approximately 18 lower at 5 nm, but reaches its saturation value for thicknesses above 20 nm, indicating cobalt interdiffusion at the silicon interfaces. The magnetization of the cobalt thicker films was the expected for an atomic magnetic moment of Co of 1.7 . The magnetization of the thickest NdCo films was approximately 800 emu/cm3. The expected magnetization of the NdCo4.6 compound, , can be calculated from the atomic concentration, the neodymium and cobalt atomic volumes determined from their related densities (8.9 gr/cm3 and 6.9 gr/cm3), and their magnetic moments and . For the atomic concentration of NdCo4.6, . The expected magnetic moment of the 4f orbital in Nd is related to the electronic configuration . However, this value is difficult to reach at RT because of the weak exchange coupling between Co and Nd atoms, disorder and possible presence of Nd unbonded to Co [22, 23]. In addition, cobalt reduces its magnetic moment by bonding to neodymium, which also contributes to the reduction of the atomic magnetic moment of Nd. This was observed in previous XMCD measurements[22] in samples with similar atomic concentrations. By taking into account all these factors, it is estimated that the magnetic moment of Co and Nd in these samples is roughly 1.35 and 1.1 , respectively. As it will be shown from the HAXPES data, these relatively low magnetic moments can hardly be attributed to oxygen contamination. The PMA energy in the thickest of the studied films, obtained from , is of 2.4 erg/cm3.
@@ III.2 Xray Reflectivity
Figure 2 compares the XRR measurements of the NdCo4.6 analyzed samples, except for the 5 nm thick sample, with the corresponding pure cobalt reference samples. The overlapping curves in red represent the related simulation curves. The simulations were done by calculating the Xray interferences between the reflections at the different layer interfaces using the Abeles formalism[40]. The parameters to adjust for each layer included in the simulation were its thickness, complex refractive index (which is related to the electron density of the material) and roughness of the interface with the adjacent top layer. The high frequency oscillations (Kiessig oscillations) in the reflectivity curves are related to the thickness of the thicker layer. The low frequency modulation in the amplitude of these oscillations corresponds to thinner layers which in this case are the silicon capping layer plus any additional layer related to segregated material. This low-frequency modulation is different in NdCo4.6 thin films compared to that of pure cobalt.
The best fits for the reflectivity curves of the cobalt films consist of a single layer of cobalt plus a capping layer of approximately 1.1 nm thickness with an electron density slightly higher than that of silicon. The exception to this fit is the 50 nm thick thin film which requires a top layer of 2.2 nm, which is slightly less dense than that of silicon, and a 1.2 nm thick interdiffusion layer.
@@
The second layer in addition to the capping top layer might be associated with segregated neodymium because this layer does not appear in the pure cobalt films. However, the electron density obtained from the best reflectivity simulations was slightly higher than half of the expected value for pure Nd. Attempts to fit the curves using metal neodymium density yielded worse fits and thinner layers. This additional layer should then be oxidized or silicized neodymium to match the observed electron densities. A possible limitation of XRR in this case is the closeness of the refraction index for cobalt and neodymium to the photon energy used, which would make Nd segregation less evident if it was mixed with cobalt.
@@
The differences in the amplitude of the Kiessig oscillations between the two orientations arise from variations in the roughness at the substrate interface. This asymmetry is also observed in thinner films, though the effect is less significant. It originates directly from the oblique incidence of Nd atoms during thin-film growth, which establishes a preferential diffusion direction perpendicular to the Nd incidence plane. As a result, the NdCo concentration becomes more homogeneous along this perpendicular direction.
This mechanism explains the in-plane magnetic anisotropy in these films as a form of shape anisotropy, driven by the reduced magnetization and weaker magnetic exchange of Nd with Co compared to Co–Co interactions at RT. However, the large energy difference between the in-plane and out-of-plane anisotropies, of one order of magnitude even in thicker samples where shadowing effects are more evident, indicates that perpendicular magnetic anisotropy (PMA) cannot be attributed solely to such inhomogeneities. Instead, the Nd environments responsible for PMA must be located in Co-rich regions, where neodymium atoms are more strongly exchange-coupled to cobalt. This interpretation is consistent with the observed increase in PMA as the films are cooled to lower temperatures [22, 14, 23].

### v4

@@ Abstract
Magnetic anisotropy in disordered rare-earth-transition metals (RE-TM) compounds arises from RE atoms occupying asymmetric environments within the TM lattice. However, the underlying mechanism that promotes such environments remains not fully understood. In this study, we investigate amorphous NdCo4.6 thin films deposited by magnetron sputtering, where the magnetic anisotropy evolves with thickness from in-plane to out-of-plane orientation above 40 nm. X-ray reflectivity measurements revealed the progressive formation of an additional layer between the 3 nm Si capping layer and the NdCo film with increasing film thickness. To probe the composition and distribution of Co and Nd near the surface, Hard X-ray Photoemission Spectroscopy (HAXPES) was performed on films ranging from 5 nm to 65 nm in thickness using incident photon energies of 7, 19, and 13 keV. These correspond to inelastic electron mean free paths of 7.2-12.3 nm in cobalt. The cobalt atomic concentration, deduced from HAXPES at the Nd 3d and Co 2p excitations, was consistently below the nominal value and varied with both thickness and photon energy. This indicates segregation of RE atoms at the film surface, which becomes more pronounced with increasing thickness. Background analysis of the Co 2p and Nd 3d peaks supports this conclusion. The thickness of the Nd segregated layer is estimated in 2-3 nm. Nd segregation provides evidence for Nd diffusion toward the surface during thin film growth, what demonstrates that incorporating neodymium into the cobalt lattice incurs an energy cost which is associated with strain due to the volume mismatch between the two elements. Reducing this strain energy promotes anisotropic atomic environments for those Nd trapped into the cobalt lattice, thereby explaining the emergence of perpendicular anisotropy and its dependence on film thickness in these RE-TM compounds.
@@ I INTRODUCTION
The PMA of RE-TM compounds is caused by the strong spin-orbit coupled 4f electron of the RE atoms. This orbital is nearly isolated from the valence atoms. Magnetic coupling between the spin of the TM and the 4f electrons occurs owing to the intratomic exchange with their valence 6s and 5d electrons which are spin polarized by the 3d TM electrons [7]. This coupling is weak, in the range of meV, and competes in energy with the crystal field at the RE[8, 7, 9] which fixes the orientation of its 4f angular moment that defines the anisotropy of the compound. The asymmetry of the RE atomic environment is, therefore, key for understanding the PMA properties of RE-TM alloys[10, 11, 12]. Structural asymmetry has been demonstrated in other RE-TM compounds such as TbFe by difraction[12], transmission electron microscopy[13] and extended Xray absorption fine spectroscopy (EXAFS)[14, 15].
Understanding how RE enters this type of asymmetric environment is an old problem owing to the complexity of the system[16, 17, 18]. However, the most accepted explanation for thin films deposited using magnetron sputtering methods, which have incident atom energies above those of thermal evaporation [19], is based on lowering the surface energy during thin film growth[20, 21]. Following this explanation, a majority of RE are in asymmetric environments because the corresponding adatom locations during growth are those that reduce the surface energy the most, creating a structural texture that gives rise to the resulting anisotropic environments. Thus, to control the PMA and the magnetic properties of the NdCo compounds it is neccesary to elucidate the parameters that condition the adatom diffusion during thin film growth.
@@
Previous XMCD measurements of the studied NdCo films performed at temperatures ranged from 10 K up to RT demonstrated the out-of-plane magnetic anisotropy of Nd and the in-plane anisotropy of Co, identifying Nd as the origin of the PMA in the alloy [24], and consequently proving that a majority of Nd atoms have to be in asymmetric atomic environments. EXAFS spectroscopy performed in 100 nm thick samples was not able to identify them due to a vanish EXAFS signal for neodymium[15], likely due to disorder [31] and an apparent low Nd concentration in the bulk. However, the measurements evidenced a highly compressed cobalt sublattice. Cobalt atoms presented a short-range order equivalent to a fcc lattice with a first neighbor distance of 2.40 Å. This is equivalent to a reduction in volume of 11 with respect to fcc cobalt at ambient pressure [32]. This bond length reduction is not present in pure cobalt films deposited in the same magnetron sputtering system or in other amorphous alloys like Co-Si at similar concentrations [33], indicating that introducing Nd into the cobalt lattice causes a significant compressive strain. This short-range ordered cobalt environment was not comparable to the expected in crystalline NdCo5 which have 6 first nearest neighbors at 2.45 Å. It also indicates larger than expected distances between Nd atoms within the bulk of the alloy. The measurements done at 10K, where thermal expansion of the film and the substrate is reduced with respect to RT, reported an anomalous expansion of the cobalt lattice, what confirmed its compressive strained structure.
@@
The present experiment shows that the PMA properties of our NdCo4.6 are thickness dependent. PMA appears above a threshold thickness between 30 to 40 nm and gradually increases until it reaches its saturation value at a thickness below 100 nm, indicating that it is built during the growth process of the film, what is coherent with an increasing compressive strain as thickness increases. In order to obtain information on the possible diffusion process occurring during thin film growth that drive the system to its PMA properties, X-ray reflectivity (XRR) and HAXPES[34, 34] were applied to different thin-film thicknesses. Both techniques are not invasive and they are sensitive to the bulk and the interfaces of the deposited films. XRR is highly sensitive to layer interfaces and should detect changes in thickness, electron density and roughness. However, it cannot provide the precise nature of the elemental composition of the layers and interfaces without ambiguity. The electron densities of cobalt and neodymium may not be large enough to be distinguished if there is some intermixing. Photoemission spectroscopy is an elemental-sensitive probe with probe depths that can be about three times the inelastic electron mean free path () of the photoemitted electrons[35]. By using hard Xrays, HAXPES reaches larger probing depths than Xray absorption spectroscopy, when using secondary electron detection [36], or Xray photoemission (XPS). Moreover, the probe depth can be modified by the incident photon energy and/or by choosing the specific excitation peaks[34]. This allows the study of the composition and chemistry of the top layer region of the samples underneath the 3 nm thick silicon capping layer and checks the extension of the possible segregated layers by changing the probe depth.
@@
The experiment is especially centered in the information collected by HAXPES to determine the gradation in the composition of thin films[37]. This information comes from the measured atomic relative concentrations of cobalt to neodymium, the valence band spectra, the shape of the photoemitted Co 2p and Nd 3d peaks and the shape and intensity of their background due to inelastically scattered electrons[38, 39]. The results of the experiment are consistent with an effective segregated layer of neodymium in all the thin films analyzed whose effective thickness is of the order of 2 to 3 nm, demonstrating that there is a diffusion of neodymium atoms toward the surface. This is consequent with the compressive strain caused by the incorporation of neodymium in cobalt. This also indicates that Nd atoms trapped within the cobalt lattice must adopt atomic environments that, in average, should reduce the compressive strain energy of the cobalt lattice. This means that they should be more relaxed in the vertical direction where lattice expansion is energetically favorable, explaining the origin of their PMA properties.
@@ II EXPERIMENT
The samples were co-deposited using three confocal magnetron sputtering guns, Co, Nd and Si. The Co gun was normal to the substrate whereas the Nd and Si guns formed an angle of 30∘ on opposite sides to the normal of the sample. The Ar pressure during deposition was 3 mbar, with a base pressure of 1 mbar. The thin films were deposited on the native oxide of boron doped silicon substrates at a rate of 2 Åsec. The voltage bias and power applied to the magnetron guns during deposition were approximately 430 V and 0.10 kW for Co, and 330 V and 0.03 kW for Nd. All the films had a protective capping layer of 3 nm thick silicon. Deposition rates and concentration were calibrated using a quartz balanced and XRR.
@@ III.1 Magnetic Characterization
Figure 1 shows the in-plane field dependent magnetization (HLs) for pure cobalt and NdCo4.6 thin films of similar thickness obtained by magnetometry at RT. The shape of the HLs of cobalt was independent of the film thickness. This is shown in figure 1(a) where the HLs of the cobalt thin films with thicknesses ranging from 5 to 50 nm are superimposed. Their coercive field is 20 Oe, which is practically the same as that measured in the NdCo4.6 thinner films (fig. 1(b)). Owing to the normal incidence of Co atoms, they have negligible in-plane magnetic anisotropy. Their magnetic remanence was close to 60.
@@
The HLs measured in the NdCo4.6 samples with thicknesses below 40 nm are typical of soft magnets with in-plane anisotropy. Figure 1(b) shows them superimposed because they barely change with thickness. Their magnetic easy axis (EA) is perpendicular to the plane of incidence of the Nd and Si atoms, confirming the oblique incidence origin of their in-plane anisotropy[42, 43, 44]. Their EA HLs have almost 100 remanence, indicating that PMA energy is small compared with the in-plane anisotropy at these thicknesses. The hard axis (HA) has a saturation field, , of approximately 200 Oe at 30 nm, increasing to almost 400 Oe at lower thicknesses. The opening of the hysteresis loop in the HA might indicate a possible out of plane anisotropy component. As the thickness of the samples increased above 40 nm, the increased, the remanence decreased and the in-plane increased, as shown in figure 1(c). These are all signs of increasing magnetic anisotropy in the alloy. The HL of the 65 nm thick sample, in figure 1(d), shows a which is 10 times the measured in the samples with thickness lower than 40 nm. In addition, there is almost no difference between the HL measured along the in-plane EA and the HA: the in-plane anisotropy becomes overwhelmed by the out-of-plane anisotropy.
@@
The magnetization of both cobalt and NdCo4.6 thin films is approximately 18 lower at 5 nm, but reaches its saturation value for thicknesses above 20 nm, indicating cobalt interdiffusion at the silicon interfaces. XRR measurements shows interfaces widths of the order of 4 to 5 Å. The dominant in-plane magnetic anisotropy at low thickness, where the magnetization is the lowest, proofs that the PMA energy at low thickness is smaller than at higher thickness. The magnetization of the cobalt thicker films was the expected for an atomic magnetic moment of Co of 1.7 . The magnetization of the thickest NdCo films was approximately 800 emu/cm3. The expected magnetization of the NdCo4.6 compound, , can be calculated from the atomic concentration, the neodymium and cobalt atomic volumes determined from their related densities (8.9 gr/cm3 and 6.9 gr/cm3), and their magnetic moments and . For the atomic concentration of NdCo4.6, . The expected magnetic moment of the 4f orbital in Nd is related to the electronic configuration . However, this value is difficult to reach at RT because of the weak exchange coupling between Co and Nd atoms, disorder and possible presence of Nd unbonded to Co [24, 25]. In addition, cobalt reduces its magnetic moment not only by bonding to neodymium but because of the reduced cobalt bond length registered by EXAFS [15, 45], which contributes to the reduction of the atomic magnetic moment of Nd. This was observed in previous XMCD measurements[24] in samples with similar atomic concentrations. By taking into account all these factors, it is estimated that the magnetic moment of Co and Nd in these samples is roughly 1.35 and 1.1 , respectively. As it will be shown from the HAXPES data, these relatively low magnetic moments can hardly be attributed to oxygen contamination.
The anisotropy energy of the analyzed samples can be estimated from . PMA energy is 0.84 erg/cm3 in the 45 nm film and 2.4 erg/cm3 in the 65 nm thick film. The in-plane anisotropy energy of the samples with 30 nm thickness or lower is approximately 0.3 erg/cm3. The apparent sharp transition from in-plane to out-of-plane anisotropy, which happens in a thickness difference of only 10 nm, is explained because the magnetic anisotropy is measured with the applied magnetic field in the plane of the samples. The transition occurs for two reasons: the gradual increase in PMA energy with thickness and the increase in thickness itself. At a critical thickness, both quantities become large enough to allow the formation of stripe domains. After this transition, figure 1 shows that the PMA energy increases gradually with thickness. The analysis of thicker samples shows that the saturation value for the PMA energy is reached around 65 nm.
@@ III.2 Xray Reflectivity
Figure 2 compares the XRR measurements of the NdCo4.6 analyzed samples, except for the 5 nm thick sample, with the corresponding pure cobalt reference samples. The overlapping curves in red represent the related fitting curves. They have larger misfits in the thinner samples indicating a higher incertitude for the parameters obtained for those films. The fitting of the reflectivity curves were done by calculating the Xray interferences between the reflections at the different layer interfaces using the Abeles formalism[46]. The parameters to adjust for each layer included in the simulation were its thickness, complex refractive index (which is related to the electron density of the material) and roughness/diffusion of the interface with the adjacent top layer. A first guess of the thickness components present in the curves was obtained from their Fourier transform. The high frequency oscillations (Kiessig oscillations) in the reflectivity curves are mainly related to the thickness of the thicker layer. The low frequency modulation in the amplitude of these oscillations corresponds to thinner layers which in this case are the silicon capping layer plus any additional layer related to segregated material. This low-frequency modulation is different in NdCo4.6 thin films compared to that of pure cobalt. The best fits for the reflectivity curves of the cobalt films consist of a single layer of cobalt plus a capping layer of approximately 1.1 nm thickness with an electron density slightly higher than that of silicon. The exception to this fit is the 50 nm thick thin film which requires a top layer of 2.2 nm, which is slightly less dense than that of silicon, and a 1.2 nm thick interdiffusion layer.
@@
The second layer in addition to the capping top layer might be associated with segregated neodymium because this layer does not appear in the pure cobalt films. However, the electron density obtained from the best reflectivity simulations was slightly higher than half of the expected value for pure Nd. Attempts to fit the curves using metal neodymium density yielded worse fits and thinner layers. This additional layer should then be oxidized or silicized neodymium to match the observed electron densities. A better identification of the elements segregated at the top layers is done by HAXPES.
@@
The differences in the amplitude of the Kiessig oscillations between the two orientations arise from variations in the roughness at the substrate interface. This asymmetry is also observed in thinner films, though the effect is less significant. It originates directly from the oblique incidence of Nd atoms during thin-film growth, which establishes a preferential diffusion direction perpendicular to the Nd incidence plane. As a result, the deposited tend to expand on the surface in the direction perpendicular to the plane of incidence creating a corrugation in the roughness of the interfaces which is anisotropic.
This mechanism explains the in-plane magnetic anisotropy in these films as a form of shape anisotropy [44]. This anisotropy might be also driven by the reduced magnetization and weaker magnetic exchange of Nd with Co compared to Co–Co interactions at RT if the distribution of neodymium is not completely homogeneous, i.e., if there is an in-plane asymmetry for the Nd environments [47]. However, the large energy difference between the in-plane and out-of-plane anisotropies, of one order of magnitude even in thicker samples where shadowing effects are more evident, indicates that the perpendicular magnetic anisotropy (PMA) is of a different nature than the in-plane one. Also, it cannot be attributed to effects like columnar shape anisotropy as occurs in other TM amorphous alloys with PMA [48]. Instead, the Nd environments responsible for PMA are those where neodymium atoms are more strongly exchange-coupled to cobalt. This interpretation is consistent with the observed increase in magnetization and PMA energy as the films are cooled to lower temperatures [24, 15, 25].